OpenStudy (anonymous):

A train is moving with a speed of 100km/h due south and a car is moving at 80km/h from the east. Find the velocity of the car relative to the train? help

OpenStudy (gorv):



OpenStudy (anonymous):

then?

OpenStudy (gorv):



OpenStudy (gorv):

got any idea...for train cae will move along resultant

OpenStudy (gorv):

u know how to find resultant of vector

OpenStudy (anonymous):

is it resultant?

OpenStudy (gorv):

yeah

OpenStudy (gorv):

can u do ???

OpenStudy (anonymous):

velocity of train 0 -100 velocity of car -80 0 then velocity of car relative to train is Vc-Vt ????

OpenStudy (gorv):

resultant vector =\[\sqrt{a^2+b^2-2a*b*\cos \theta}\]

OpenStudy (anonymous):

(-80 (0 -80 0) MINUS -100) = 100

OpenStudy (anonymous):

ITS THE ANSWER

OpenStudy (anonymous):

??

OpenStudy (gorv):

what i dint get what u type ??